OpenStudy (anonymous):

How do you find the anti derivative of (1-|x|)?

myininaya (myininaya):

if x>0, then |x|=x if x<0, then |x|=-x hope that helps

OpenStudy (anonymous):

Maybe I am not asking the question right. its the intrgural from -1 to 1\[\int\limits_{-1}^{1}(1-|X|)dx\]

OpenStudy (anonymous):

Sorry does that make more sense?

myininaya (myininaya):

\[\int\limits_{-1}^{1}|x|dx=\int\limits_{-1}^{0}(-x) dx+\int\limits_{0}^{1}(x) dx\]

myininaya (myininaya):

\[\frac{-x^2}{2}|_{-1}^{0}+\frac{x^2}{2}|_0^1=[0-\frac{-1}{2}]+[\frac{1}{2}-0]=1\]

myininaya (myininaya):

so \[\int\limits_{-1}^{1}(1-|x|)dx=\int\limits_{-1}^{1}1 dx-\int\limits_{-1}^{1}|x| dx\] \[=x|^1_{-1}-1=[1-(-1)]-1=1+1-1=2-1=1\]